OpenStudy (anonymous):

Which expression is equivalent to 12x βˆ’ 8x + 3 and why? 3 + 4x, because (12x βˆ’ 8x) + 3 = 4x + 3 = 3 + 4x 7x, because 12x βˆ’ 8x + 3 = 4x + 3 = 7x 12x βˆ’ 5, because 12x βˆ’ (8x βˆ’ 3) = 12x βˆ’ 5 x, because 12x βˆ’ 8x + 3 = 12x βˆ’ 11x = x

OpenStudy (anonymous):

3 + 4x, because (12x βˆ’ 8x) + 3 = 4x + 3 = 3 + 4x